OpenStudy (anonymous):

4 C3H5N3O9---> 6 N2+ 12 CO2 + 10 H2O + O2 This reaction generates a large amount of heat and many gaseous products. It is the sudden formation of these gases, together with a rapid expansion, that produces the explosion. a) what is the maximum amount of O2 in grams that can be obtained from 2.00x10^2 of nitroglycerin? b) calculate the percent yield in this reaction if the same amount of O2 generated is found to be 5.87g.

OpenStudy (anonymous):

ok first of is that 2*10^2 moles of nitro? if yes than your answer is 1600 g of O2 can be produced by folowing calculation: n(o2)/n(nitro) = 1/4 --> n(O2) = 1/4 * 2*10^2 = 50 mol n=m/M --> m= n*M m(O2) = 50 mol * 32 gmol-1 = 1600 g and seccond part with yeald is it question if from that same amount (2*10^2 mol) you get 5,87 g of O2? if so than it is: yeald = 5,87 g/ 1600 g = 3,67*10^-3 *100 = 0,367%